• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Zolnierczuk

    Log entry time 09:07:42 on August22,2001

    Entry number 71835

    keyword=target EPR=42.2%/NMR=40.9%

    8:04 lasers #1,2,4 off
    target moved to pickup coil position
    8:05 Rotate field from 270 -> 0deg (BL=-7.2042A,BT=-0.1179A)
    8:10 NMR 40.87%
    8:13 lasers #5,6,7 on -> 8:17
    8:22 EPR begin (RTD#6 187.0 / RTD#7 191.5)
    trouble finding signal - it turned out that the EPR FG was left on 0.85Vpp
    although it read 0.025. The Vpp button was not hit after latest EPR
    average P=42.2%
    8:39 target moved to BeO, 0deg, lasers 5,6,7 on


    EPR fit details
    0th order fit
    freq1: -47.75 kHz +/- 3.79 kHz | Polarisation: 44.66 % +/- 3.55 %
    freq2: 43.02 kHz +/- 4.28 kHz | Polarisation: 40.23 % +/- 4 %

    1st order fit
    freq1: -45.7 kHz +/- 3.55 kHz | Polarisation: 42.74 % +/- 3.32 %
    freq2: 43.91 kHz +/- 4.14 kHz | Polarisation: 41.06 % +/- 3.87 %